Find the equation of the plane through the line of intersection of the planes x-z = 3 and y+3z = 4 and perpendicular to the plane x+y+z = 1.
Answer #1

ANSWER:-

Given that

the planes x-z = 3 and y+3z = 4

and perpendicular to the plane x+y+z = 1.

The lanes - Z=3 and y + 32=4 . perpendicular to the plane at yt2=1 Equation of a plane which passes through intersection Pinosubstitute ge the value & in eq 0 to the equation of Plane. x-olyst. (0-1) Z- (3 + 4 (0)=0 - 1-2-3-0.